BabyGurl2006:

How do i do slope intercept form...an example of a problem im stuck on: Write an equation in slope-intercept form for the line with slope -1/4 and y-intercept 2.

bri05:

do u know what the formula is to write it?

BabyGurl2006:

no, i forget

bri05:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @BabyGurl2006 no, i forget \(\color{#0cbb34}{\text{End of Quote}}\) ok its not that hard give me a min

bri05:

ok so the formula to write it out is y=mx+b, now m is the slope and b is the y-intercept. Now that u now that what do you think that answer to ur question is

BabyGurl2006:

ummmm... y=-1/4x+2 ???

bri05:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @BabyGurl2006 ummmm... y=-1/4x+2 ??? \(\color{#0cbb34}{\text{End of Quote}}\) yes u got it

BabyGurl2006:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @bri05 \(\color{#0cbb34}{\text{Originally Posted by}}\) @BabyGurl2006 ummmm... y=-1/4x+2 ??? \(\color{#0cbb34}{\text{End of Quote}}\) yes u got it \(\color{#0cbb34}{\text{End of Quote}}\) oh wow...thnx!

bri05:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @BabyGurl2006 \(\color{#0cbb34}{\text{Originally Posted by}}\) @bri05 \(\color{#0cbb34}{\text{Originally Posted by}}\) @BabyGurl2006 ummmm... y=-1/4x+2 ??? \(\color{#0cbb34}{\text{End of Quote}}\) yes u got it \(\color{#0cbb34}{\text{End of Quote}}\) oh wow...thnx! \(\color{#0cbb34}{\text{End of Quote}}\) mhm ofc its not that hard u just have to remember the formula is all

pretty123:

do yall know this A store purchases an item wholesale for $60 and marks it up by 70 percent. If there is a 7 percent tax on the item, how much will the item cost?